Baker Green's weekly gross earnings for the week ending December 7 were $2,500.00, and her federal income tax withholding was $525.00. Social security rate is 6.2% and Medicare is 1.45%, and all earnings are subject to FICA taxes. She contributes a fixed amount of $75.00 per pay check to her retirement account and also has $25.00 withdrawn, towards a charitable contribution. What is Baker Green's net pay? Round your answer to two decimal places. Show your work. (10 points)
Answer #1

As mentioned in question:

Gross earnings = $2500

Federal income tax = 525

Retirement contributions = $75

Charitable contributions = $25

Social Security tax = Gross earnings * social security rate = ($2500*6.2%) = 155

Medicare tax = Gross earnings * medicare tax rate = ($2500*1.5%) = 37.5

Greens Net Pay = Gross earnings - Federal income tax -  Retirement and Charitable contributions - Social Security tax -Medicare tax

Green's Net pay = 2500 - 525 - 75 - 25 - 155 - 37.50 = $1682.50